# SAMBO-RL: Shifts-aware Model-based Offline Reinforcement Learning

### v2

@@ 4.2 Shifts-aware Reward via Probabilistic Inference
This subsection introduces a novel shifts-aware reward through a unified probabilistic inference framework, serving as a lower-bound surrogate for the true objective. We examine how this modified reward contributes to refining the value function and motivating $\pi$ to escape the limitations of behavior policy, thus mitigating model bias and policy shift.
@@
Shifts-aware Reward. We introduce the shifts-aware reward $\tilde{r}(s_{t},a_{t},s_{t+1})$ formulated within trajectories. It comprises the vanilla reward, adjustments for model bias, and modifications for policy shift, expressed as the following:
$\begin{split}&\tilde{r}(s_{t},a_{t},s_{t+1})=\log r(s_{t},a_{t})\\ &+\frac{1}{(1-\gamma)\gamma^{t}}\left[\log\frac{p(s_{t+1}|s_{t},a_{t})}{q(s_{t+1}|s_{t},a_{t})}+\log\frac{\pi(a_{t}|s_{t})}{\pi_{c}(a_{t}|s_{t})}\right].\end{split}$ (5)
@@
This modified reward forms the basis of our surrogate objective function during training, which is formulated as:
@@
Additionally, we establish that $\mathcal{L}(\pi)$ serves as a lower bound for the true objective, ensuring its rationality as a surrogate.
Theorem 4.1.
@@
Optimization objective based on this shifts-aware reward provides a methodological foundation for integrating shift adjustments seamlessly into reinforcement learning algorithms to uniformly address model bias and policy shift.
Modal Bias Adjustment. We now examine how this adjustment refines the value estimation and addresses the challenge of model bias. By rolling out the learned policy $\pi$ within the model dynamics $m$, we obtain a model-generated dataset $\mathcal{D}_{m}$. Furthermore, the modified training objective of the model dataset $\mathcal{D}_{m}$ can be computed as the following:
@@
This modified reward consists of the logarithmic vanilla reward penalized by the KL divergence between the model and environment dynamics. This adjustment reduces value function misestimations by accounting for the accuracy of the learned model, encouraging the development of both a high-return policy and an accurate model approximation.
@@
Policy Shift Modification. To delve into the positive effect of policy shift modification, we consider a dataset $\mathcal{D}_{b}$ generated by executing the behavior policy $\pi_{b}$ within the environment $p$. Moreover, the modified training objective of the dataset $\mathcal{D}_{b}$ is derived as the following formulation:
$\tilde{J}(\pi)=\mathbb{E}_{s\sim\mathcal{D}_{b},a\sim\pi(\cdot|s),s^{\prime}\sim\mathcal{D}_{b}}\left[\tilde{r}(s,a,s^{\prime})\right]$
$=\mathbb{E}_{s\sim\mathcal{D}_{b}}\left[\mathbb{E}_{a\sim\pi(\cdot|s)}\left[\log r(s,a)\right]+\beta D_{\operatorname{KL}}(\pi(\cdot|s)||\pi_{b}(\cdot|s))\right],$
which introduces a bonus to the vanilla reward based on the KL divergence between $\pi$ and $\pi_{b}$. This modification guides $\pi$ to break out of the confines of the behavior policy $\pi_{b}$, promoting better exploration and more effective learning.
4.3 Shifts-aware Model-based Offline RL
@@
We apply the shifts-aware reward in model-based offline RL and develop the practical framework, SAMBO-RL.
@@
To adapt the shifts-aware reward for practical implementation, where training data is typically stored in transition format rather than trajectories, we introduce two weighting factors, $\alpha$ and $\beta$. These factors replace the respective coefficients in the reward function, shaping it as the following:
$\tilde{r}(s,a,s^{\prime})=\log r(s,a)+\alpha\log\frac{p(s^{\prime}|s,a)}{q(s^{\prime}|s,a)}+\beta\log\frac{\pi(a|s)}{\pi_{c}(a|s)}.$
@@
These weighting factors ensure the reward and adjustments are properly balanced in transition-based learning scenarios, which is empirically validated in the following experiments.
@@
Furthermore, since the environment dynamics $p$ and behavior policy $\pi_{b}$ are typically unknown and cannot be exactly computed, we derive two classifiers to estimate the shifts-aware reward (Eysenbach et al., 2020, 2022). Details are shown in Appendix A.3. Specifically, these classifiers estimate model bias adjustment $\log\frac{p(s^{\prime}|s,a)}{q(s^{\prime}|s,a)}$ and policy shift modification $\log\frac{\pi(a|s)}{\pi_{c}(a|s)}$ respectively. The transition classifier $C_{\phi}(s,a,s^{\prime})$ estimates the probability that the transition $(s,a,s^{\prime})$ comes from the environment, while the action classifier $C_{\psi}(s,a)$ estimate the probability that a given state-action pair $(s,a)$ is generated by the current policy $\pi$.
We train the classifiers by minimizing the loss functions:
@@
For the transition classifier $C_{\varphi}=C_{\phi}(s,a,s^{\prime})$, $\mathcal{D}_{1}=\mathcal{D}_{env}$ is the offline dataset, and $\mathcal{D}_{2}=\mathcal{D}_{m}$ is the model-generated dataset. For the action classifier $C_{\varphi}=C_{\psi}(s,a)$, $\mathcal{D}_{1}=\mathcal{D}_{\pi}$ is the dataset collected by executing the current policy $\pi$, and $\mathcal{D}_{2}=\mathcal{D}_{env}$ is the offline dataset. Based on these classifiers, we approximate the shifts-aware reward as the following:
$\tilde{r}(s,a,s^{\prime})\approx\log r(s,a)$ (8)
@@
$+\alpha\log\frac{C_{\phi}(s,a,s^{\prime})}{1-C_{\phi}(s,a,s^{\prime})}+\beta\log\frac{C_{\psi}(s,a)}{1-C_{\psi}(s,a)}.$
@@
The complete procedure of SAMBO-RL is outlined in Algorithm 1, with further details provided in Appendix B.
5 Experiments
@@
In our experiments, we focus on three objectives: (1) validating the effectiveness of shifts-aware reward, (2) comparing the performance of SAMBO with previous methods, and (3) analyzing the contribution of each component and hyperparameter within SAMBO to its overall performance.
@@
We evaluate our algorithm using standard offline RL benchmarks in Mujoco environments and conduct an ablation study to assess its performance. Our implementation is based on the OfflineRL-Kit library [^1] https://github.com/yihaosun1124/OfflineRL-Kit, a comprehensive and high-performance library for implementing offline reinforcement learning algorithms. The basic parameters of our algorithm are consistent with the settings in this library.
5.1 Effectiveness of Shifts-aware Reward
(a) Model Bias Adjustment
(b) Policy Shift Modification
Figure 2: Assessment of the shifts-aware reward. (a) Model bias adjustment mitigates inaccuracies in value estimation, achieving a nearly optimal return. (b) Policy shift modification addresses implicit policy restrictions, leading to faster convergence.
We conduct an evaluation of the effectiveness of the shifts-aware Reward in addressing model bias and policy shift issues within the 1D grid world illustrated in Fig. 1 (a).
@@ 5.1 Effectiveness of Shifts-aware Reward
Table 1: Results on the D4RL Gym benchmark. MOPO’s results are obtained on the “v0” datasets; MOPO*’s results are obtained from experiments in the OfflineRL-Kit library on the “v2” datasets. The numbers reported for SAMBO are the normalized scores averaged over the final iteration of training across 4 seeds, with $\pm$ standard deviation. The top three scores are highlighted in bold. We use an asterisk “*” for the best score, an underscore “_” for the second, and bold text for the third.
@@
Task Name BC CQL TD3+BC EDAC MOPO MOPO* COMBO TT RAMBO MOBILE SAMBO (Ours)
halfcheetah-random 2.2 31.3 11.0 28.4 35.4 38.5 38.8 6.1 39.5 39.3 39.7$\pm$2.0
hopper-random 3.7 5.3 8.5 25.3 11.7 31.7 17.9 6.9 25.4 31.9 32.4$\pm$0.5
@@
halfcheetah-medium 43.2 46.9 48.3 65.9 42.3 72.4 54.2 46.9 77.9 74.6 72.5$\pm$3.8
hopper-medium 54.1 61.9 59.3 101.6 28.0 62.8 97.2 67.1 87.0 106.6 99.7$\pm$0.3
walker2d-medium 70.9 79.5 83.7 92.5 17.8 84.1 81.9 81.3 84.9 87.7 78.3$\pm$2.6
halfcheetah-medium-replay 37.6 45.3 44.6 61.3 53.1 72.1 55.1 44.1 68.7 71.7 70.6$\pm$1.6
hopper-medium-replay 16.6 86.3 60.9 110.0 67.5 92.7 89.5 99.4 99.5 103.9 101.5$\pm$1.5
walker2d-medium-replay 20.3 76.8 81.8 87.1 39.0 85.9 56.0 82.6 89.2 89.9 92.3$\pm$ 1.5
halfcheetah-medium-expert 44.0 95.0 90.0 106.3 63.3 83.6 90.0 95.0 95.4 108.2 97.8 $\pm$0.3
hopper-medium-expert 53.9 96.9 98.0 110.7 23.7 74.6 111.0 110.0 88.2 112.6 111.0 $\pm$1.2
@@
Model Bias Adjustment. Model bias exacerbates inaccuracies in value estimates, ultimately resulting in a distorted optimization objective. To evaluate the effects of model bias adjustment on the training process, we conduct experiments. As illustrated in Fig. 2 (a), we execute the policy gradient algorithm under four settings: using the vanilla reward with the overestimating model (OM) or underestimating model (UM), the shifts-aware reward with the overestimating model (SAR with OM) underestimating model (SAR with UM). We then evaluate the expected return of the learned policy after each training update in the actual environment. The results show that while the OM initially achieves optimal returns, it rapidly diverges from the optimal policy. In contrast, SAR with OM effectively counters the performance degradation caused by model overestimation, sustaining optimal returns. Similarly, SAR with UM mitigates the policy suboptimality due to model underestimation, achieving a return near the theoretical optimum.
(a) MOBILE
(b) SAMBO
@@
Figure 3: The $(s,a,s^{\prime})$ distributions of offline RL algorithms result in the Hopper-M task. Model Data represents the distribution of data generated by executing the final learned policy in the dynamic model, while Real Data represents the distribution of data generated by executing the final learned policy in the environment.
Policy Shift Modification. Policy shift implicitly restricts policy updates. To assess the efficacy of the policy shift modification, we apply the policy gradient algorithm to the offline objective $\mathcal{J}(\pi)=\mathbb{E}_{s\sim d^{p,\pi_{b}}}\left[V^{\pi}(s)\right]$ in the environment depicted in Fig. 1 (a). We test four settings: the vanilla reward or SAR, combined with a behavior policy that is either uniformly distributed or the opposite of the optimal policy. We measure the KL divergence between the learned policy and the behavior policy after each update. As illustrated in Fig. 2 (b), the algorithm using the vanilla reward is limited by the behavior policy, serving a lower KL divergence and slower convergence. In contrast, the algorithm employing the SAR effectively motivates policy to be optimistic. It overcomes the restriction, displaying a higher and more stable divergence and achieving faster convergence.

### v3

@@ 4.2. Shifts-aware Reward via Probabilistic Inference
This subsection introduces a novel shifts-aware reward through a unified probabilistic inference framework, serving as a lower-bound surrogate for the true objective. We examine how this modified reward contributes to refining the value function and constraining unstable policy updates, thereby mitigating the distribution shift in both training and real-world deployment.
@@
Shifts-aware Reward. We introduce the shifts-aware reward $\tilde{r}(s_{t},a_{t},s_{t+1})$ formulated within trajectories. It comprises the vanilla reward, adjustments for model bias and modifications for policy shift, expressed as the following equation:
$\begin{split}\tilde{r}(s_{t},a_{t},s_{t+1})&=\log r(s_{t},a_{t})\\ &+\frac{1}{(1-\gamma)\gamma^{t}}\left[\log\frac{p(s_{t+1}|s_{t},a_{t})}{q(s_{t+1}|s_{t},a_{t})}+\log\frac{\pi(a_{t}|s_{t})}{\pi_{c}(a_{t}|s_{t})}\right].\end{split}$ (5)
@@
This modified reward forms the basis of our surrogate objective function during training, which is defined as:
@@
Additionally, we establish that the surrogate objective $\mathcal{L}(\pi)$ serves as a lower bound for the true objective.
Theorem. 1.
@@
The theorem ensures that $\mathcal{L}(\pi)$ is a rational surrogate objective for $\mathcal{J}_{M}(\pi)$. Optimization objective based on this shifts-aware reward provides a methodological foundation for integrating shift adjustments seamlessly to uniformly address distribution shift caused by model bias and policy shift.
Model Bias Adjustment. We now examine how this adjustment refines the value estimation and addresses the challenge of model bias. By rolling out the learned policy $\pi$ in model $m$, we obtain a model-generated dataset $\mathcal{D}_{m}$. Then, when training on dataset $\mathcal{D}_{train}$, the modified objective restricted to $\mathcal{D}_{m}$ is computed as:
@@
This modified reward consists of the logarithmic vanilla reward penalized by the KL divergence between the model and environment dynamics. This adjustment reduces value function misestimation by accounting for the accuracy of the learned model. By emphasizing data from regions with reliable model predictions, it mitigates the distribution shift in training data caused by model bias.
@@
Policy Shift Modification. To delve into the positive effect of policy shift modification, we consider the modified training objective in the training dataset $\mathcal{D}_{train}$ restricted to $\mathcal{D}_{env}$, in which case the dynamic $q$ satisfies $q=p$ and the data collection policy satisfies $\pi_{c}=\pi_{b}$ in equation (4). Then the modified objective is
$\tilde{J}(\pi)$ $=\mathbb{E}_{(s,a,s^{\prime})\sim\mathcal{D}_{env}}\left[\tilde{r}(s,a,s^{\prime})\right]$
$=\mathbb{E}_{s\sim\mathcal{D}_{env}}\left[\mathbb{E}_{a\sim\pi_{b}(\cdot|s)}\left[\log r(s,a)\right]-\beta D_{\operatorname{KL}}(\pi_{b}(\cdot|s)||\pi(\cdot|s))\right].$
By introducing a penalty to the vanilla reward based on the KL divergence between offline data collection policy $\pi_{b}$ and testing policy $\pi$, this modification prevents $\pi$ from making overly aggressive updates, thus stabilizing the learning process. It reduces the bias between the offline data and the testing policy, helping to mitigate the distribution shift caused by policy shift.
4.3. Shifts-aware Model-based Offline RL
@@
We apply the shifts-aware reward in model-based offline RL and develop the algorithm SAMBO-RL.
@@
To adapt the shifts-aware reward for practical implementation, where training data is typically stored in transition format rather than trajectories, we introduce two weighting factors, $\alpha$ and $\beta$. These factors replace respective coefficients in reward function (5):
$\tilde{r}(s,a,s^{\prime})=\log r(s,a)+\alpha\log\frac{p(s^{\prime}|s,a)}{q(s^{\prime}|s,a)}+\beta\log\frac{\pi(a|s)}{\pi_{c}(a|s)}.$ (7)
@@
These weighting factors ensure the reward and adjustments are properly balanced in transition-based learning scenarios, which is empirically validated in the following experiments. Moreover, as discussed in Section 4.2, for data in $\mathcal{D}_{env}$, the dynamics satisfy $q=p$ and the policy satisfies $\pi_{c}=\pi_{b}$. For data in $\mathcal{D}_{m}$, we have $q=m$, and the data-collection policy can be approximated by $\pi$, i.e., $\pi_{c}=\pi$. As discussed in detail in Appendix B.2, we have
$\tilde{r}(s,a,s^{\prime})=\begin{cases}\log r(s,a)+\beta\log\frac{\pi(a|s)}{\pi_{b}(a|s)},&\forall(s,a,s^{\prime})\in\mathcal{D}_{env}\\ \log r(s,a)+\alpha\log\frac{p(s^{\prime}|s,a)}{m(s^{\prime}|s,a)},&\forall(s,a,s^{\prime})\in\mathcal{D}_{m}\end{cases}$
@@
Furthermore, since the environment dynamics $p$ and the policy $\pi_{b}$ are typically unknown and cannot be exactly computed, we derive classifiers to estimate the shifts-aware reward Eysenbach et al. (2020, 2022); Liu et al. (2021). Details are shown in Appendix A.2. Specifically, these classifiers estimate $\log\frac{p(s^{\prime}|s,a)}{m(s^{\prime}|s,a)}$ and $\log\frac{\pi(a|s)}{\pi_{b}(a|s)}$ respectively. The transition classifier $C_{\phi}(s,a,s^{\prime})$ estimates the probability that the transition $(s,a,s^{\prime})$ comes from the environment, while the action classifier $C_{\psi}(s,a)$ estimate the probability that a given state-action pair $(s,a)$ is generated by the current policy $\pi$. We train these classifiers by minimizing the cross-entropy loss:
@@
For the transition classifier $C_{\varphi}=C_{\phi}(s,a,s^{\prime})$, $\mathcal{D}_{1}=\mathcal{D}_{env}$ and $\mathcal{D}_{2}=\mathcal{D}_{m}$. For the action classifier $C_{\varphi}=C_{\psi}(s,a)$, $\mathcal{D}_{1}=\mathcal{D}_{\pi}$ is the dataset collected by executing the current policy $\pi$, and $\mathcal{D}_{2}=\mathcal{D}_{env}$. Based on these classifiers, we approximate the shift-aware reward by
@@
$\tilde{r}(s,a,s^{\prime})=\begin{cases}\log r(s,a)+\beta\log\frac{C_{\psi}(s,a)}{1-C_{\psi}(s,a)},&\forall(s,a,s^{\prime})\in\mathcal{D}_{env}\\ \log r(s,a)+\alpha\log\frac{C_{\phi}(s,a,s^{\prime})}{1-C_{\phi}(s,a,s^{\prime})},&\forall(s,a,s^{\prime})\in\mathcal{D}_{m}\end{cases}$ (9)
@@
The procedure of SAMBO-RL is outlined in Algorithm 1, with implementation details provided in Appendix B.
Algorithm 1 SAMBO-RL
1:Offline dataset $\mathcal{D}_{env}$, learned dyanmics models $\{m_{\theta}^{i}\}_{i=1}^{N}$, initialized policy $\pi$, hyperparameters $\alpha$ and $\beta$.
2:Train the probabilistic dynamics model $m_{\theta}(s^{\prime},r|s,a)=\mathcal{N}(\mu_{\theta}(s,a),\Sigma_{\theta}(s,a))$ on $\mathcal{D}_{env}$.
3:Initialize the model dataset buffer $\mathcal{D}_{m}\leftarrow\varnothing$.
4:for $i=1,2,...,N_{iter}$ do
5:  Initialize the policy dataset buffer $\mathcal{D}_{\pi}\leftarrow\varnothing$.
6:  Generate synthetic rollouts by model $m_{\theta}$. Add transition data in these rollouts to $\mathcal{D}_{m}$ and $\mathcal{D}_{\pi}$.
7:  Update classifiers $C_{\phi}$ and $C_{\psi}$ according to Eq. (8).
8:  Compute shifts-aware reward according to Eq. (9).
9:  Run SAC Haarnoja et al. (2018) with shifts-aware reward to update policy $\pi$.
10:end for
4.4. Theoretical Analysis
We analyze the theoretical properties of SAR, with detailed proofs provided in Appendix A. We first introduce $\xi$-uncertainty quantifier, which characterizes the suboptimality of the derived policy.
Definition 2.
($\xi$-Uncertainty Quantifier Jin et al. (2021)). The set of penalization $\{\Gamma_{h}\}_{h\in[H]}$ forms a $\xi$-uncertainty quantifier if it holds with probability at least $1-\xi$ that
$\left|\widehat{\mathbb{B}}\widehat{V}_{t+1}(s,a)-\mathbb{B}\widehat{V}_{t+1}(s,a)\right|\leq\Gamma_{t}(s,a),\quad\forall(s,a)\in\mathcal{S}\times\mathcal{A}$
where $\widehat{V}_{t+1}$ is an estimated value function at step $t+1$ and $\widehat{\mathbb{B}}$ is an empirical Bellman operator estimating the Bellman operator $\mathbb{B}$.
We can reformulate (7) in a state-action style:
$\tilde{r}(s,a)$ $=\mathbb{E}_{s^{\prime}\sim q(\cdot|s)}\left[\tilde{r}(s,a,s^{\prime})\right]$
$=\log r(s,a)-\alpha D_{\operatorname{KL}}(q(\cdot|s,a)\|p(\cdot|s,a))-\beta\log\frac{\pi_{c}(a|s)}{\pi(a|s)}.$
Since the model cannot perfectly fit the environment dynamics $p$, we assume that the KL divergence satisfies $D_{\operatorname{KL}}(m(\cdot\mid s,a)\,\|\,p(\cdot\mid s,a))\geq\epsilon_{KL}>0$. Under this KL divergence assumption, we prove that the penalty of SAR $\tilde{r}(s,a)$ forms a $\xi$-uncertainty quantifier.
Theorem. 3.
Under the assumption that $D_{\operatorname{KL}}(m(\cdot\mid s,a)\,\|\,p(\cdot\mid s,a))\geq\epsilon_{KL}>0$, there exist $\alpha>0$ and $\beta>0$ such that the penalty of SAR forms a $\xi$-uncertainty quantifier.
Based on this theorem, we derive the following value bound from the perspective of pessimistic value iteration (PEVI) Jin et al. (2021), which demonstrates that SAR is an effective corrective reward.
Theorem. 4.
Let $\pi^{*}$ be the optimal policy in the environment, $\pi^{*}_{G}$ be the optimal policy in MDP $G=(\mathcal{S},\mathcal{A},p,\log r,\mu_{0},\gamma)$, $V_{G}^{\pi}(s)$ the value function in $G$ and $\epsilon_{s}\geq D_{\operatorname{KL}}(m(\cdot\mid s,a)\,\|\,p(\cdot\mid s,a))\geq\epsilon_{KL}>0$. Then the derived policy $\widehat{\pi}$ according to SAR satisfies
$\left|V^{\pi^{*}}(s_{1})-V^{\widehat{\pi}}(s_{1})\right|\leq 2\alpha H\epsilon_{s}+\epsilon(G;s_{1}),$
with probability at least $1-\xi$ for all $s\in\mathcal{S}$. Here,
$\epsilon(G;s)=|V^{\widehat{\pi}}(s)+V_{G}^{\pi^{*}_{G}}(s)-V^{\pi^{*}}(s)-V_{G}^{\widehat{\pi}}(s)|,$
which quantify the value error between different MDPs and $\epsilon(M;s)=0,\forall s\in\mathcal{S}$.
5. Experiments
@@
In our experiments, we focus on three objectives: (1) validating the effectiveness of shifts-aware reward, (2) comparing the performance of SAMBO with previous model-free and model-based methods, and (3) analyzing the contribution of each component in SAMBO.
@@
We evaluate our algorithm using standard offline RL benchmarks in Mujoco environments and conduct an ablation study to assess its performance. Our implementation is based on the OfflineRL-Kit library [^1] https://github.com/yihaosun1124/OfflineRL-Kit, a comprehensive and high-performance library for implementing offline RL algorithms. The basic parameters of our algorithm are consistent with the settings in this library.
@@
Table 1. Results on the D4RL Gym benchmark. MOPO are obtained on the “v0” datasets; MOPO* are obtained from experiments in the OfflineRL-Kit library on the “v2” datasets. The numbers reported for SAMBO are the normalized scores averaged over the final iteration of training across 4 seeds, with $\pm$ standard deviation. The top three scores are highlighted in bold.
@@
Task Name BC CQL TD3+BC EDAC MOPO MOPO* COMBO $T$ RAMBO MOBILE(SOTA) SAMBO (Ours)
[…]
@@
[…]
halfcheetah-medium-expert 44.0 95.0 90.7 106.3 63.3 83.6 90.0 95.0 95.4 108.2 97.8 $\pm$0.3
hopper-medium-expert 53.9 96.9 98.0 110.7 23.7 74.6 111.1 110.0 88.2 112.6 111.0 $\pm$1.2
@@
[…]
(a) SAR in Toy Example
(b) MOBILE
(c) SAMBO
@@
Figure 2. Effectiveness of the shifts-aware reward. (a) SAR is capable of achieving a near-optimal return. (b) and (c) The $(s,a,s^{\prime})$ distributions of offline RL algorithms result in the Hopper-M task in NeoRL benchmark. Model Data represents the distribution of data generated by executing the final learned policy in the dynamic model, while Real Data represents the distribution of data generated by executing the final learned policy in the real environment.
5.1. Effectiveness of Shifts-aware Reward
Effectiveness in Toy Example. To assess the efficacy of shifts-aware reward in different models, we examine its amendment to the training process. As illustrated in Fig. 2 (a), we execute the policy gradient algorithm under four settings: using the vanilla reward with the overestimating model (OM) or underestimating model (UM), the SAR with the overestimating model (SAR with OM) or underestimating model (SAR with UM). We then evaluate the expected return of the learned policy after each training update in the real environment. The results show that while the OM initially achieves optimal returns, it rapidly diverges from the optimal policy. In contrast, SAR with OM effectively counters the performance degradation caused by model overestimation, sustaining optimal returns. Similarly, SAR with UM mitigates the policy suboptimality due to model underestimation, achieving a return near the theoretical optimum. Both demonstrate the effectiveness.